OpenStudy (anonymous):

Which of the following ordered pairs is a solution to the system of linear equations below? 3x+y=9} 2x - 2y=14} A (2, 3) B (3, 2) C (4,-3) D (-3,4)

OpenStudy (anonymous):

My bad just dont forget to look at the right of your screen please and thank you

OpenStudy (anonymous):

C is a solution to both. 3(4)+(-3)=9 and 2(4)-2(-3)=14

hero (hero):

Solve for y for both equations, set y = y, then solve for x. Input x into the original equation, then solve for y.

OpenStudy (anonymous):

the question basically says that both lines intersect at one of the following coordinates, and which one is it? therfore, one of the coordinates in the agrees with both linear equations

OpenStudy (anonymous):

in the choices*

hero (hero):

Working on it...stand by

hero (hero):

Whoops made a mistake...recalculating: Solve both equations for y: y = (14 - 2x)/-2 y = (9-3x) Set y = y: (14-2x)/-2 = (9-3x) Solve for x: -2(9-3x) = 14-2x -18 + 6x = 14 - 2x 8x = 32 x = 4 y = -3